OpenStudy (nwolf):

I need HELP!! what is the solution of the system 6x-9y=25.4 7y-2x=10.5

OpenStudy (mysesshou):

6x-9y=25.4 7y-2x=10.5 By addition method, 6x - 9y = 25.4 ( -2x + 7y = 10.5 ) * 3 gives 6x - 9y = 25.4 -6x + 21y = 31.5 add them. 6x - 9y = 25.4 -6x + 21y = 31.5 = 12y = 56.9 y= 4.74 plug y into one of the original equations to solve for x. x=? OR......... a different method to solve it.... by substitution method: 6x-9y=25.4 (eq1) 7y-2x=10.5 (eq2) 7y-2x=10.5 (eq2) solve for y 7y=10.5+2x y = (2/7)x + (10.5/7) substitute y into the other equation (the first one, since this y is from the second one) 6x-9y=25.4 (eq1) 6x-9[ (2/7)x + (10.5/7) ]=25.4 solve for x: 6x-[ (2*9/7)x + (10.5*9/7) ]=25.4 6x-[ (2.57)x + (13.5) ]=25.4 6x-(2.57)x - (13.5) =25.4 3.43x = 38.9 x = 11.34 then substitute x into one of the original equations and solve for the y. but, for grins and giggles, let's check my x and y 6(11.34)-9(4.74)=25.4 6(11.34)-9(4.74)=25.4 68.04 - 42.66 = 25.4 25.38 = 25.4 <-- works, but is off slightly due to rounding. but if you round the left side, you do get 25.4 = 25.4, so the values for x and y are right. hope this explains some. This was 2 different methods you could use.